# Global Dynamics of McKean-Vlasov SDEs via Stochastic Order

### v1

@@ Acknowledgement
We acknowledge the financial supports of EPSRC grant ref (ref. EP/S005293/2), Royal Society through the Award of Newton International Fellowship (ref. NIF/R1/221003) and National Natural Science Foundation of China (No. 12171280).
@@ Appendix A Nonexistence of Strong Ordered Pairs for Stochastic Order
In addition to closedness of the stochastic order mentioned in Section 3.1, almost all classical results in the theory of systems that preserve a certain partial order (monotone dynamical systems) are established on the fundamental assumption that the state space $X$ is not only endowed with a partial order relation $\leq$”, but is also a “strongly ordered space” (see e.g., Hirsch [41, 42]). A strongly ordered space implies that the interior of the partial order relation $\leq$” (referred to as the strong ordering $\ll$”) is a nonempty relation on $X$, and its closure is $\leq$”. The strong ordering $\ll$” is a relation for which two points “$x\ll y$” means that there are open neighbourhoods $U,V$ in $X$ of $x,y$ respectively such that $U\leq V$, where the term “$U\leq V$” refers to $u\leq v$ for any $u\in U$ and $v\in V$. If $x\ll y$, we call $(x,y)$ is a strong ordered pair. Based on the tool of order topology, especially the topology generated by the interiors of order intervals enclosed by strong ordered pairs, the theory of monotone dynamical systems and its applications to cooperative irreducible ODEs (coordinate order in $\mathbb{R}^{d}$), delay equations and parabolic equations (pointwise order in functional spaces) have undergone extensively investigation (see e.g., [60, 43, 71]).
Unfortunately, in the appendix, we will show that the strong ordering relation generated by the stochastic order “$\leq_{\mathrm{st}}$” in $\mathcal{P}_{2}(\mathbb{R})$ is empty, meaning that NO strong ordered pairs exist. We actually prove a stronger conclusion – for any $\mu\in\mathcal{P}_{2}(\mathbb{R})$, every open set $B\subset\mathcal{P}_{2}(\mathbb{R})$ has an element that is order unrelated with $\mu$ (see Proposition A.1). This implies that classical theory of monotone dynamical systems are inapplicable to order-preserving semigroups on $\mathcal{P}_{2}(\mathbb{R})$. Consequently, exploration of properties of the stochastic order and establishment of the theory for systems preserving the stochastic order, as discussed in Section 3.2, 4.1 and 5.3, are indispensable.
@@
[…]
@@
[…]
If $\mu$ and $\nu_{1}$ are order unrelated, the proof is completed by choosing $\nu=\nu_{1}$. Otherwise, without loss of generality, we assume $\mu\leq_{\mathrm{st}}\nu_{1}$. Then we know that
@@
[…]
@@
Hence, it follows from Theorem 6.9 in [68] again that $\lim_{n\to\infty}\mathcal{W}_{2}(\nu_{2,n},\nu_{1})=0.$ Thus, there exists $n_{0}\geq c$ such that $\nu:=\nu_{2,n_{0}}\in B(\nu_{1},\delta)\subset B(\nu_{0},2\delta)$. By (A.5), the construction of $\nu_{2,n_{0}}$ in (A.6) and the choice of $r_{0}$ in Case 1, we know that
@@
Similarly, we can conclude that $\nu_{3,n}\in\mathcal{P}_{2}(\mathbb{R})$ and $\lim_{n\to\infty}\mathcal{W}_{2}(\nu_{3,n},\nu_{1})=0.$ Choose $n_{0}\geq c$ large enough such that $\nu:=\nu_{3,n_{0}}\in B(\nu_{1},\delta)\subset B(\nu_{0},2\delta)$, then we know that for all $r\geq r_{n_{0}}$,

### v2

@@ Acknowledgement
We acknowledge the financial supports of EPSRC grant ref (ref. EP/S005293/2), Royal Society through the Award of Newton International Fellowship (ref. NIF/R1/221003) and National Natural Science Foundation of China (No. 12501184, No. 12171280).
@@ Appendix A Nonexistence of Strong Ordered Pairs for Stochastic Order
In addition to closedness of the stochastic order, almost all classical results in the theory of monotone dynamical systems are established on the fundamental assumption that the state space $X$ is not only endowed with a partial order relation $\leq$”, but is also a “strongly ordered space” (see e.g., Hirsch [27, 28]). A strongly ordered space implies that the interior of the partial order relation $\leq$” (referred to as the strong ordering $\ll$”) is a nonempty relation on $X$, and its closure is $\leq$”. The strong ordering $\ll$” is a relation for which two points $x\ll y$” means that there are open neighbourhoods $U,V$ in $X$ of $x,y$ respectively such that $U\leq V$, where the term $U\leq V$” refers to $u\leq v$ for any $u\in U$ and $v\in V$. If $x\ll y$, we call $(x,y)$ is a strong ordered pair. Based on the tool of order topology, especially the topology generated by the interiors of order intervals enclosed by strong ordered pairs, the theory of monotone dynamical systems and its applications to cooperative irreducible ODEs, delay equations and parabolic equations have undergone extensively investigation (see e.g., [47, 26, 59]).
Unfortunately, in the appendix, we will show that the strong ordering relation generated by the stochastic order $\leq_{\mathrm{st}}$” in $\mathcal{P}_{2}(\mathbb{R})$ is empty, meaning that NO strong ordered pairs exist. We actually prove a stronger conclusion – for any $\mu\in\mathcal{P}_{2}(\mathbb{R})$, every open set $B\subset\mathcal{P}_{2}(\mathbb{R})$ has an element that is order unrelated with $\mu$ (see Proposition A.1). This implies that classical theory of monotone dynamical systems are inapplicable to order-preserving semigroups on $\mathcal{P}_{2}(\mathbb{R})$. Consequently, exploration of properties of the stochastic order and the results of monotone dynamical systems, as discussed in Section 2.2, 3.1 and 4.2, are indispensable.
@@
where $|f|_{\infty}=\sup_{x\in\mathbb{R}}|f(x)|$. Hence $\nu_{1,\epsilon}$ weakly converges to $\nu_{0}$ as $\epsilon\to 0$, which is denoted by $\nu_{1,\epsilon}\xrightarrow{w}\nu_{0}$ as $\epsilon\to 0$. Moreover, similar to (A.2), we have
@@
By Theorem 6.9 in [54], we have $\lim_{\epsilon\to 0}\mathcal{W}_{2}(\nu_{1,\epsilon},\nu_{0})=0$ due to $\nu_{1,\varepsilon}\xrightarrow{w}\nu_{0}$ and (A.3). Then there exists $0<\epsilon_{1}<\epsilon_{0}$ such that $\nu_{1}:=\nu_{1,\epsilon_{1}}\in B(\nu_{0},\delta)$ satisfying (A.1). Thus, we have proved the Claim.
If $\mu$ and $\nu_{1}$ are order unrelated, the proof is completed by choosing $\nu=\nu_{1}$. Otherwise, without loss of generality, we assume $\mu\leq_{\mathrm{st}}\nu_{1}$. Then we have
@@
It is easy to check that $\nu_{2,n}\in\mathcal{P}_{2}(\mathbb{R})$ for all $n\geq c$. Moreover, $\nu_{2,n}\xrightarrow{w}\nu_{1}$ as $n\to\infty$ since for any $f\in C_{b}(\mathbb{R})$,
@@
Hence, it follows from Theorem 6.9 in [54] again that $\lim_{n\to\infty}\mathcal{W}_{2}(\nu_{2,n},\nu_{1})=0.$ Thus, there exists $n_{0}\geq c$ such that $\nu:=\nu_{2,n_{0}}\in B(\nu_{1},\delta)\subset B(\nu_{0},2\delta)$. By (A.5), the construction of $\nu_{2,n_{0}}$ in (A.6) and the choice of $r_{0}$ in Case 1, we have
@@
Similarly, we can conclude that $\nu_{3,n}\in\mathcal{P}_{2}(\mathbb{R})$ and $\lim_{n\to\infty}\mathcal{W}_{2}(\nu_{3,n},\nu_{1})=0.$ Choose $n_{0}\geq c$ large enough such that $\nu:=\nu_{3,n_{0}}\in B(\nu_{1},\delta)\subset B(\nu_{0},2\delta)$, then we have for all $r\geq r_{n_{0}}$,